OpenStudy (anonymous):

ASAP MEDAL! solve the equation sqrt11x-5=sqrt9x+7

OpenStudy (tkhunny):

Is this it? \(\sqrt{11x-5} = \sqrt{9x+7}\) Have you first considered the Domain?

OpenStudy (anonymous):

yes, and no. can you also please explain step by step as to how to get the answer? I would like to understand it.

OpenStudy (tkhunny):

You must consider the Domain. \(11x-5 \ge 0\) Do you know why? \(9x+7 \ge 0\) Do you know why?

OpenStudy (anonymous):

i do not. can you explain

OpenStudy (tkhunny):

Can you find the square root of a negative number?

OpenStudy (anonymous):

honestly im so bad at math... im sorry I sound super dumb it;s just not my strong subject..

OpenStudy (tkhunny):

Never say that again. Now, answer my question. \(\sqrt{4} = 2\) because 2*2 = 4 What is \(\sqrt{-4}\)? Can you multiply two identical REAL numbers and get -4?

OpenStudy (anonymous):

-2*-2

OpenStudy (tkhunny):

Does that work? (-2)(-2) = +4 We did NOT get -4. Try again.

OpenStudy (anonymous):

oh no then? cause -2*2=-4 but they aren't the same?

OpenStudy (tkhunny):

Perfect!!! If we are going to stay in the Real Numbers, the is NO Square root of nagative numbers. Keep this important fact in mind. We have \(\sqrt{11x-5}\), so 11x - 5 BETTER NOT BE negative. Agreed?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Then, \(11x - 5 \ge 0\) or \(11x\ge 5\) or \(x \ge 5/11\). You do the other side. \(\sqrt{9x+7}\)

OpenStudy (anonymous):

Im so confused as to how you did that...

OpenStudy (tkhunny):

What is it that needs not to be negative?

OpenStudy (anonymous):

so 9x+7>=0 or 9x>=7 or x>=7/9?

OpenStudy (tkhunny):

-7/9 Do you see the error?

OpenStudy (tkhunny):

The second to last thing we need is a little thing called equivalence. \(If \sqrt{a} = \sqrt{b}\;then\;a = b\). Some might construe this as Squaring Both Sides, but that idea leads to more errors than it should. We have \(\sqrt{11x-5} = \sqrt{9x+7}\), so 11x-5 = 9x+7. Solve for x and make sure the result is in the established Domain, x >= -7/9